Sunset rents an SUV at $21.95 plus $0.23 per mile. Sunrise rents the same vehicle for $24.95 plus $0.19 per mile. For what mileage is the cost the same? Hint: Using "m" for miles and "c" for cost, write one equation for each rental company. Then solve the system.
Answer by checkley71(8403)

WHY USE 2 EQUATIONS WHEN 1 WILL DO AS FOLLOWS:
21.95+.23M=24.95+.19M
.23M-.19M=24.95-21.95
.04M=3
N=3/.04
M=75 MILES IS THE BREAK EVEN FIGURE FR THESE 2 OFFERS.
PROOF"
21.95+.23*75=24.95+.19*75
21.95+17.25=24.95+14.25
39.20=39.20